Solve the following differential equations:
i) (x²+y²)dy/dx=xy
1
Expert's answer
2019-04-28T17:35:02-0400

Solution. We write the equation as


"(x^2+y^2)dy=xydx"

"-xydx+(x^2+y^2)dy=0"

Consider the equation as


"M(x,y)dx+N(x,y)dy=0."

Find


"\frac {\delta M} {\delta y} = - x"

"\frac {\delta N} {\delta x} = 2x""\frac {\delta M} {\delta y} =\not \frac {\delta N} {\delta x}"

Find the value


"-\frac {\frac {\delta M} {\delta y}-\frac {\delta N} {\delta x}} {M(x,y)} =-\frac {-x-2x} {-xy} =- \frac {3} {y}."

The resulting function depends only on y. Therefore, find the integrating factor using the equation


"\frac {k'(y)} {k(y)} = -\frac {3} {y}."

Hence


"k(y)=\frac {1} {y^3}"

.

Get the total (exact) differential equation


"-\frac {x} {y^2}dx+(\frac {x^2} {y^3}+\frac {1} {y})dy=0."

Let a solution to the equation be U(x,y). Hence


"\frac {\delta U} {\delta x} = -\frac {x} {y^2}"

Get


"U=- \frac {x^2} {2y^2} +C(y)"

where C(y) is function of y. Therefore


"\frac {\delta U} {\delta y} =\frac {x^2} {y^3}+C'(y)=\frac {x^2} {y^3}+\frac {1} {y}"

"C'(y)=\frac {1} {y}""\u0421(y)=ln(y) + C"

where C is constant. Hence


"U(x,y)=- \frac {x^2} {2y^2}+ln(y)+C=k,"

where k is constant.

Answer: 

"U(x,y)=- \frac {x^2} {2y^2}+ln(y)=D"

where D is constant.